# Fast and Accurate Algorithms to Calculate Expected Modularity in Probabilistic Networks

### v2

@@ 3.1 Possible world partitioning for expected modularity
(b) A tree enumerating all partitions of possible worlds defined by community $c$.
(c) All possible worlds in partition 7.
Figure 2: An example demonstrating how, for a given community, all possible worlds of a probabilistic graph can be partitioned so that $|e_{c}^{w}|,|e_{c,\bar{c}}^{w}|$, and $|e_{\bar{c}}^{w}|$ are constant inside each partition.
3.2 Approximate partitioning for expected modularity
In Eq. 15, $F_{\varrho}$ will contain $\frac{T_{\varrho}!}{(T_{\varrho}-\varrho)!\varrho!}$ elements, the sum over which is infeasible to compute in practice. Approximate Possible World Partitioning for Expected Modularity ($\mathrm{APWP}^{\mathrm{EMOD}}$) is an approximation variation of $\mathrm{PWP}^{\mathrm{EMOD}}$ where we use Discrete Fourier Transform (DFT) to approximate the probabilities for disjoint sets $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$ and $e_{\bar{c}}^{w}$ [23, 24],
$\begin{split}Pr(|e_{c}^{w}|=x)&\approx\frac{1}{T_{x}+1}\sum_{l=0}^{T_{x}}C^{-lx}\prod_{\alpha=1}^{T_{x}}(1+(C^{l}-1)p_{\alpha}^{x}),\end{split}$ (16)
$\begin{split}Pr(|e_{c,\bar{c}}^{w}|=y)&\approx\frac{1}{T_{y}+1}\sum_{l=0}^{T_{y}}C^{-ly}\prod_{\alpha=1}^{T_{y}}(1+(C^{l}-1)p_{\alpha}^{y}),\end{split}$ (17)
$\begin{split}Pr(|e_{\bar{c}}|=z)&\approx\frac{1}{T_{z}+1}\sum_{l=0}^{T_{z}}C^{-lz}\prod_{\alpha=1}^{T_{z}}(1+(C^{l}-1)p_{\alpha}^{z}).\end{split}$ (18)
The probability of partition $xyz$ is then:
$\begin{split}Pr(xyz)&\approx\prod_{\varrho\in\{x,y,z\}}\frac{1}{T_{\varrho}+1}\sum_{l=0}^{T_{\varrho}}C^{-l\varrho}\prod_{\alpha=1}^{T_{\varrho}}(1+(C^{l}-1)p_{\alpha}^{\varrho}),\end{split}$ (19)
where $C=exp(\frac{2i\pi}{T_{\varrho}+1})$, $i=\sqrt{-1}$, and $p_{\alpha}^{x}$, $p_{\alpha}^{y}$, $p_{\alpha}^{z}$ are, respectively, the probabilities of edge $\alpha$ in $e_{c}^{w},e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$.
By substituting Eq. 19 into Eq. 9, we obtain,
$\begin{split}E(\textbf{Q})&=\sum_{c\in\mathcal{C}}\sum_{x=1}^{T_{x}}\sum_{y=1}^{T_{y}}\sum_{z=1}^{T_{z}}Q_{c}^{xyz}Pr(xyz)\\ &\approx\sum_{c\in\mathcal{C}}\sum_{x=1}^{T_{x}}\sum_{y=1}^{T_{y}}\sum_{z=1}^{T_{z}}[\frac{x}{(x+y+z)}-(\frac{2x+y}{2(x+y+z)})^{2}]\\ &\times\prod_{\varrho\in\{x,y,z\}}\frac{1}{T_{\varrho}+1}\sum_{l=0}^{T_{\varrho}}C^{-l\varrho}\prod_{\alpha=1}^{T_{\varrho}}(1+(C^{l}-1)p_{\alpha}^{\varrho}).\end{split}$ (20)
@@ 3.2 Approximate partitioning for expected modularity
$\mathrm{APWP}^{\mathrm{EMOD}}$ is also detailed in Algorithm 2. In this case, differently from $\mathrm{PWP}^{\mathrm{EMOD}}$, in lines 6, 8 and 10 we use Eq. 16, Eq. 17 and Eq. 18 to calculate $Pr(|e_{c}^{w}|=x)$, $Pr(|e_{c,\bar{c}}^{w}|=y)$ and $Pr(|e_{\bar{c}}^{w}|=z)$ separately.
@@
output : Approximated expected modularity: $Sum$
@@ 17                         $Sum=Sum+Q_{c}^{xyz}P_{c}[x]P_{c,\bar{c}}[y]P_{\bar{c}}[z]$
Algorithm 2 $\mathrm{(A)PWP}^{\mathrm{EMOD}}$
Computational complexity: The worst-case time complexity of $\mathrm{APWP}^{\mathrm{EMOD}}$ is $O(k|E|^{3})$. We can break the analysis down into two main parts.
Lines 3 to 10 concern the approximation of probabilities for each partition. The time complexities of $Pr(x)$, $Pr(y)$, and $Pr(z)$ are $O(|E|^{2})$. From Equation 9, we can see that $x$ ranges from 1 to $T_{x}$, $y$ ranges from 1 to $T_{y}$, and $z$ ranges from 1 to $T_{z}$. $T_{x}$, $T_{y}$, and $T_{z}$ sum to $|E|$, so the overall complexity is $O(|E|^{3})$.
@@
Lines 11 to 15 involve looping over all partitions, which has a time complexity of $O(|E|^{3})$. Notice that in the worst-case scenario, the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ are similar to each other. Here the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ have the same upper bound $|E|$.
@@
Both parts are inside the community loop on line 2, so the time complexity of the whole algorithm is $O(k|E|^{3})$, where $k$ is the number of communities. Notice that the time complexity solely depends on the number of communities and edges, and remains unaffected by the assignment of edge probabilities.
4 Evaluation
@@
This section has three main objectives. First, we show that our algorithms are as accurate as the brute-force method , but are several orders of magnitude faster. Notice that the the brute-force method iterates over all possible worlds and thus gives us the correct result. Then, we use our approximation algorithm to evaluate the main current alternatives to estimate expected modularity: weighting, thresholding, and sampling. Finally, we study the behaviour of our algorithm when we vary the number of communities, the distribution of community sizes, and the type of input network.
@@
All the experiments except the comparison with the brute-force method have been performed on a macOS system, with max CPU frequency 2.4GHz. The comparison with the brute-force method has been performed on a Linux system, with max CPU frequency 5GHz. Both systems have the same 32GB memory capacity. All the code is in Python and available at https://github.com/uuinfolab/Expected Modularity Calculation in Probabilistic Graph.
@@
Datasets. In our experiments we use both synthetic datasets, to control and examine the properties of the data that may affect running time and accuracy of the tested algorithms, and real-world datasets, to demonstrate the practical use of our methods.
@@
To generate synthetic datasets, we use Stochastic Block Model, Forest Fire Network (FFN), Barabási-Abert (BA), Small world (SW), and Erdős-Rényi (ER). The parameters used to generate specific random graphs are specified later in this section for each experiment.
@@
As real datasets, we use the Enron email network, a Facebook network, a protein-protein interaction (PPI) network, and a collaboration network, summarized in Table 2. The data sources used to generate these networks have been commonly used in the literature on probabilistic networks. In particular, two datasets (Enron, PPI) have been directly taken from the literature, while for those not provided by the authors we obtained probabilistic networks from the same sources following similar procedures and reproducing similar network statistics. For the two datasets we reproduced and PPI dataset, to keep the original topology structure, we included all the nodes in selected clusters detected by modularity optimization and all edges between those nodes. In our experiments, these clusters are also used as input to compute expected modularity.
@@
Enron: this dataset consists of emails sent between employees of Enron between 1999 and 2001. Nodes represent employees and there is an edge between two nodes if at least one email has been exchanged between them. The original dataset with edge probabilities is provided by the authors in [7].
@@
Protein-protein interaction (PPI): the dataset contains nodes that represent the proteins, edges that represent the interactions between two proteins, and associated probabilities. This dataset is extracted from a protein database that directly provides interaction probabilities, as described in [1].
@@
Facebook: the dataset is obtained from a weighted Facebook-like Forum network, that originates from an online community for students at the University of California, Irvine. The data, downloaded from toreopsahl.com, includes the users who sent or received at least one message. Probabilities are computed using the method proposed in [25]. Specifically, we applied an exponential cumulative distribution function (CDF) of mean 2 to the weights. We sample nodes and edges from the original networks.
[…]
@@
4.1 Accuracy and execution time
To show the accuracy of our methods, we compare the values of expected modularity computed using $\mathrm{PWP}^{\mathrm{EMOD}}$ and $\mathrm{APWP}^{\mathrm{EMOD}}$ with the true results, that we can obtain using a brute-force approach iterating through all the possible worlds. It is worth noting that $\mathrm{PWP}^{\mathrm{EMOD}}$ is not an approximation algorithm, and is expected to compute the exact same result as the brute-force method: the experimental results are provided to show the correctness of the mathematical derivations on which $\mathrm{PWP}^{\mathrm{EMOD}}$ is based. On the contrary, $\mathrm{APWP}^{\mathrm{EMOD}}$ computes an approximation of expected modularity with the aim of further reducing computing time. Therefore, this experiment also aims to inspect the accuracy of the approximation.
For the first experiment, we generate networks with three communities and varying sizes, using a stochastic block model (also known as planted-community structure network) [27], with the parameters shown in Table 3. To control the size of networks, we make $p_{in}$ and $p_{out}$ sometimes slightly different. For the second experiment, we generate random probabilistic graphs, using Erdős–Rényi model, with different numbers of edges and we fix the number of communities to 5. Table 4 shows that the results calculated by $\mathrm{PWP}^{\mathrm{EMOD}}$ and $\mathrm{APWP}^{\mathrm{EMOD}}$ are always the same as the true values, calculated by the brute-force method, down to several significant digits (four in the table). We also observe that even for larger networks where we can no longer use the brute-force method, the results obtained by $\mathrm{APWP}^{\mathrm{EMOD}}$ is practically the same as the one produced by $\mathrm{PWP}^{\mathrm{EMOD}}$, which does not involve any approximations.
@@ 4.1 Accuracy and execution time
$m$ $S^{\mathrm{PWP}}$ $S^{\mathrm{APWP}}$ $S^{BF}$
@@
$S^{\mathrm{PWP}}$ expected modularity score of $\mathrm{PWP}^{\mathrm{EMOD}}$
$S^{\mathrm{APWP}}$ expected modularity score of $\mathrm{APWP}^{\mathrm{EMOD}}$
$S^{BF}$ expected modularity score of brute-force method
Table 4: Expected modularity computed using the brute-force method, $\mathrm{PWP}^{\mathrm{EMOD}}$, and $\mathrm{APWP}^{\mathrm{EMOD}}$.
Table 5 and Figure 3 show that both our methods are multiple orders of magnitude faster than the brute-force approach. Figure 3 adds larger networks to test the behavior of $\mathrm{APWP}^{\mathrm{EMOD}}$ where the two other methods cannot be used. Both $\mathrm{PWP}^{\mathrm{EMOD}}$ and the brute-force method have an exponential time complexity, although, $\mathrm{PWP}^{\mathrm{EMOD}}$ is significantly faster. Figure 3 confirms the polynomial time complexity of $\mathrm{APWP}^{\mathrm{EMOD}}$.
$m$ $T^{PWP}$ $T^{APWP}$ $T^{BF}$
@@
$T^{PWP}$ running time of $\mathrm{PWP}^{\mathrm{EMOD}}$
$T^{APWP}$ running time of $\mathrm{APWP}^{\mathrm{EMOD}}$
@@
Table 5: Running time of brute-force , $\mathrm{PWP}^{\mathrm{EMOD}}$, and $\mathrm{APWP}^{\mathrm{EMOD}}$.
Figure 3: Running time of brute-force, $\mathrm{PWP}^{\mathrm{EMOD}}$, and $\mathrm{APWP}^{\mathrm{EMOD}}$, with log axes.
In the next sections we only use $\mathrm{APWP}^{\mathrm{EMOD}}$, given that it is both accurate and much faster than the brute-force method and $\mathrm{PWP}^{\mathrm{EMOD}}$.
4.2 Comparison with alternative methods
In this section we show that both weighting (that is, treating the probabilities as weights) and thresholding (that is, considering the network as a deterministic one after keeping only high-probability edges) lead to wrong estimations of expected modularity. On the contrary, sampling can be used to get more and more accurate results by increasing execution time, so we study how the balance between time and accuracy when using a sampling approach compares to our algorithm.
4.2.1 Weighting
In this experiment, we show that directly regarding probabilistic networks as weighted deterministic networks leads to a wrong expected modularity calculation. In particular, we first generate a base network using stochastic block modeling, where the parameters are $k=3$, $\mathrm{nc}=9$, $p_{in}=0.72$ and $p_{out}=0.12$ From this base network, we generate multiple probabilistic networks by assigning probabilities to its edges. I n each probabilistic network we assign the same probability to all the edges: for example, in Figure 4, the network corresponding to the value 0.20 on the $x$ axis has all its edges having probability 0.20.
@@ 4.2.1 Weighting
Figure 4: Weighted and expected modularity calculated on the same graphs.
4.2.2 Thresholding
@@
In this experiment, we compare expected modularity (as computed by $\mathrm{APWP}^{\mathrm{EMOD}}$) with the modularity obtained after thresholding, showing that the choice of threshold largely influences the final results, which are also not accurate in general.
@@
Figure 5: Mean and expected modularity calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ and using the threshold method.
(a) CCS graph
(b) LCCS graph
(c) CCS graph
(d) LCCS graph
(e) CCS graph
(f) LCCS graph
@@
Figure 6: Running time and expected modularity value calculated by different methods for graphs with a more clear (first row) and less clear (second row) modular structure, with different entropies.
@@
Here we generate planted community structure networks with 110 edges and 3 communities. All networks have the same topology, generated by a stochastic block model with $p_{in}=0.72$ and $p_{out}=0.12$. We then generate different probabilistic networks by assigning random probabilities to the edges so that, for different networks, entropy ratio ranges from 0 to 1, and randomly assign those probabilities to the edges. For each network, we compute the modularities obtained using different thresholds (from 0.1 to 1). In Figure 5 we plot their mean (blue circles) and standard deviation (thick vertical lines). We can see that except for networks with low entropy ratios (that is, networks very close to deterministic networks), the modularities obtained through thresholding are far from the correct values (red line, calculated using $\mathrm{APWP}^{\mathrm{EMOD}}$).
4.2.3 Sampling
@@
Thanks to the fact that using $\mathrm{APWP}^{\mathrm{EMOD}}$ we can compute expected modularity in polynomial time for a given input network, we can now test the ability of the sampling method to accurately estimate expected modularity, and study the factors influencing the balance between accuracy and execution time when sampling is used.
@@
We use two kinds of random graphs, one is characterized by a pronounced community structure, we call it CCS graph, featuring dense intra-community edges and sparse inter-community edges; another exhibits a less distinct community structure, we call it LCCS graph. Both graphs contain 27 nodes and 3 communities, with similar number of edges by moving a few edges from within a community to between community from CCS graph to LCCS graph. The CCS graph and LCCS graph are constructed with stochastic block models where the parameters are, respectively, $k=3$, $\mathrm{nc}=9$, $p_{in}=0.99$, and $p_{out}=0.01$, and $k=3$, $\mathrm{nc}=9$, $p_{in}=0.72$ and $p_{out}=0.12$. Notice that the CSS graph closely approximates three cliques. When the entropy ratio equals 1.00, all edge probabilities are assigned to 0.50; when the entropy ratio equals 0.47, all edge probabilities are assigned to 0.90; when the entropy ratio is 0.00, all edges are deterministic.
@@
Figure 6 shows the execution time (vertical blue line) and the computed expected modularity (blue horizontal line) for our algorithm. Each red circle is the result of a different execution of the sampling method. Specifically, the red nodes arranged in ’bands’ are the result of a different execution of the sampling method, all employing the same input $\theta$ as defined in Algorithm 1. For different executions, we have let the sampling method run for specific amounts of time (starting from the time needed by our algorithm), so that we could inspect the balance between execution time and likelihood to return an accurate result for the sampling method. For example, in Figure 6(f) we can see eight executions of the sampling method close to the blue vertical line: these were stopped after the same time used by our algorithm to return, and their $y$ value is the expected modularity they returned. From the figure we can see that, for different edge probability entropies, sampling always converges to the result produced by $\mathrm{APWP}^{\mathrm{EMOD}}$. Except for the case with no uncertainty (entropy ratio=0), sampling takes longer than $\mathrm{APWP}^{\mathrm{EMOD}}$ to converge toward the right results. By looking at Figures 6(b)6(d)6(f) or Figures 6(a)6(c)6(e), we find that when entropy ratio increases (larger uncertainty), sampling converges more slowly. If we compare two figures on the same column, e.g. 6(c) and 6(d), we find that when the graph structure is less clear, the convergence time of sampling is longer.
4.3 Factors influencing execution time

### v3

@@ 5             Draw a uniformly random number $r_{ij}\in[0,1]$;
Figure 3: A tree enumerating all partitions of possible worlds defined by community $c$.
Figure 4: All possible worlds in partition 7. $|e_{c}^{w}|,|e_{c,\bar{c}}^{w}|$, and $|e_{\bar{c}}^{w}|$ are constant inside the partition.
Fast/Fourier Possible-World Partitioning (FPWP)
While PWP practically only can be used in small networks, its partitioning approach, combined with the Poisson Binomial distribution, enables a further reduction in computational complexity. The Poisson Binomial distribution can be expressed in closed-form using the Discrete Fourier Transform [34];
$\begin{split}{{\color[rgb]{0,0,0}Pr(X=k)=\frac{1}{n+1}\sum_{l=0}^{n}C^{-lk}\prod_{i=1}^{n}(1+(C^{l}-1)p_{i}^{k})\ }}\end{split}$ (15)
where $C=exp(j\frac{2\pi}{n+1})$ and $j=\sqrt{-1}$.
This leads to a version of our method that we call Fast/Fourier Possible World Partitioning for Expected Modularity ($\mathrm{FPWP}$), replacing equations 12, 13, 14 by the following equations 16, 17, 18, respectively:
$\begin{split}Pr(|e_{c}^{w}|=x)&{{\color[rgb]{0,0,0}=}}\frac{1}{T_{x}+1}\sum_{l=0}^{T_{x}}C^{-lx}\prod_{\alpha=1}^{T_{x}}(1+(C^{l}-1)p_{\alpha}^{x})\ ,\end{split}$ (16)
$\begin{split}Pr(|e_{c,\bar{c}}^{w}|=y)&{{\color[rgb]{0,0,0}=}}\frac{1}{T_{y}+1}\sum_{l=0}^{T_{y}}C^{-ly}\prod_{\alpha=1}^{T_{y}}(1+(C^{l}-1)p_{\alpha}^{y})\ ,\end{split}$ (17)
$\begin{split}Pr(|e_{\bar{c}}|=z)&{{\color[rgb]{0,0,0}=}}\frac{1}{T_{z}+1}\sum_{l=0}^{T_{z}}C^{-lz}\prod_{\alpha=1}^{T_{z}}(1+(C^{l}-1)p_{\alpha}^{z})\ .\end{split}$ (18)
@@
$\mathrm{FPWP}$ is also detailed in Algorithm 2. In this case, differently from $\mathrm{PWP}$, in lines 6, 8 and 10 we use Eq. 16, Eq. 17 and Eq. 18 to calculate $Pr(|e_{c}^{w}|=x)$, $Pr(|e_{c,\bar{c}}^{w}|=y)$ and $Pr(|e_{\bar{c}}^{w}|=z)$ separately.
@@
output : Expected modularity: $Sum$
@@ 17                         $Sum=Sum+Q_{c}^{xyz}P_{c}[x]P_{c,\bar{c}}[y]P_{\bar{c}}[z]$;
Algorithm 2 $\mathrm{(F)PWP}$
@@
The worst-case time complexity of $\mathrm{FPWP}$ is $O(km^{3})$. We can break the analysis down into two main parts. Lines 3 to 10 concern the probabilities for each partition. The time complexities of $Pr(x)$, $Pr(y)$, and $Pr(z)$ are $O(m^{2})$. From Equation 9, we can see that $x$ ranges from 1 to $T_{x}$, $y$ ranges from 1 to $T_{y}$, and $z$ ranges from 1 to $T_{z}$. $T_{x}$, $T_{y}$, and $T_{z}$ sum to $m$, so the overall complexity is $O(m^{3})$. Lines 11 to 15 involve looping over all partitions, which has a time complexity of $O(m^{3})$. Notice that in the worst-case scenario, the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ are similar to each other. Here the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ have the same upper bound $m$.
@@
Both parts are inside the community loop on line 2, so the time complexity of the whole algorithm is $O(km^{3})$, where $k$ is the number of communities. Notice that the time complexity solely depends on the number of communities and edges, and remains unaffected by the assignment of edge probabilities.
Evaluation
@@
The brute-force method produces the correct result and can thus be used to evaluate the other approaches. However, it can only be used on very small networks. Therefore, we start our evaluation showing that our method is as accurate as the brute-force method, but is several orders of magnitude faster. In this way, we can then use our method to evaluate the other possible ways to estimate expected modularity: weighting, thresholding, and sampling. We also study the behavior of our algorithm when we vary the number of communities, the distribution of community sizes, and the type of input network.
@@
All the experiments except the comparison with the brute-force method have been performed on a macOS system, with max CPU frequency 2.4GHz. The comparison with the brute-force method has been performed on a Linux system, with max CPU frequency 5GHz. Both systems have the same 32GB memory capacity.
@@
Data. In our experiments we use both synthetic datasets, to control and examine the properties of the data that may affect running time and accuracy of the tested algorithms, and real-world datasets.
@@
To generate synthetic datasets, we use Stochastic Block Model (SBM), Forest Fire Network (FFN), Barabási-Abert (BA), Small World (SW), and Erdős-Rényi (ER). The parameters used to generate specific random networks are specified later in this section for each experiment.
@@
As real datasets, we use the Enron email network, an Online Social Network (OSN), a Protein-Protein Interaction (PPI) network, and a Collaboration network, summarized in Table 2. The data sources used to generate these networks have been commonly used in the literature on probabilistic networks. In particular, two datasets (Enron, PPI) have been directly taken from the literature, while for those not provided by the authors we obtained probabilistic networks from the same sources following similar procedures and reproducing similar network statistics. For OSN, Collaboration, and PPI data we included all the nodes in selected clusters detected by modularity optimization and all edges between those nodes. In our experiments, these clusters are also used as input to compute expected modularity.
  1. 1.
@@
    Enron: this dataset consists of emails sent between employees of Enron between 1999 and 2001. Nodes represent employees and there is an edge between two nodes if at least one email has been exchanged between them. The original dataset with edge probabilities is provided by the authors [7].
  2. 2.
@@
    Protein-Protein Interaction (PPI): the dataset contains nodes that represent proteins, edges that represent the interactions between two proteins, and associated probabilities. This dataset is extracted from a protein database that directly provides interaction probabilities [1].
  3. 3.
@@
    Online Social Network (OSN): the dataset is obtained from a weighted Facebook-like social network, that originates from an online community for students at the University of California, Irvine. The data, downloaded from toreopsahl.com, includes the users who sent or received at least one message. Probabilities are computed applying an exponential cumulative distribution function (CDF) of mean 2 to the weights [35]. We sample nodes and edges from the original networks.
  4. 4.
    Collaboration: the dataset is obtained from a weighted network of coauthorships between scientists posting preprints of the astrophysics archive at www.arxiv.org from 1995-1999 [36]. We computed the probabilities using the same method as for the OSN data.
@@
Accuracy and execution time
To verify the accuracy of our method, we compare the values of expected modularity computed using $\mathrm{PWP}$ and $\mathrm{FPWP}$ with the correct results, that we obtain using a brute-force approach iterating through all the possible worlds. It is worth noting that both $\mathrm{PWP}$ and $\mathrm{FPWP}$ are exact algorithms, that is, they are expected to compute the same result as the brute-force method (modulo numerical approximation): the experimental results are provided to show the correctness of our mathematical derivations.
For the first experiment, we generate networks with three communities and varying sizes, using a stochastic block model (also known as planted-community structure network) [37], with the parameters shown in Table 3. We adjust $p_{in}$ and $p_{out}$ to control the size of the networks. For the second experiment, we generate random probabilistic networks, using Erdős–Rényi model, with different numbers of edges and we fix the number of communities to 5. Table 4 shows that the results calculated by $\mathrm{PWP}$ and $\mathrm{FPWP}$ are always the same as the true values, calculated by the brute-force method, down to several significant digits (four in the table). We also observe that even for larger networks where we can no longer use the brute-force method, the results obtained by $\mathrm{FPWP}$ is the same as the one produced by $\mathrm{PWP}$, as expected.
@@
$m$ $Q^{\mathrm{PWP}}$ $Q^{\mathrm{FPWP}}$ $Q^{BF}$
@@
$Q^{\mathrm{PWP}}$ expected modularity score computed by $\mathrm{PWP}$
$Q^{\mathrm{FPWP}}$ expected modularity score computed by $\mathrm{FPWP}$
$Q^{BF}$ expected modularity score computed by the brute-force method
Table 4: Expected modularity computed using the brute-force method, $\mathrm{PWP}$, and $\mathrm{FPWP}$.
Table 5 and Figure 5 show that both our methods are multiple orders of magnitude faster than the brute-force approach. Figure 5 adds larger networks to test the behavior of $\mathrm{FPWP}$ where the two other methods cannot be used. Both $\mathrm{PWP}$ and the brute-force method have an exponential time complexity, although, $\mathrm{PWP}$ is significantly faster. Figure 5 confirms the polynomial time complexity of $\mathrm{FPWP}$.
$m$ $T^{PWP}$ $T^{FPWP}$ $T^{BF}$
@@
$T^{PWP}$ running time of $\mathrm{PWP}$
$T^{FPWP}$ running time of $\mathrm{FPWP}$
@@
Table 5: Running time of brute-force, $\mathrm{PWP}$, and $\mathrm{FPWP}$.
Figure 5: Running time of brute-force, $\mathrm{PWP}$, and $\mathrm{FPWP}$, with log axes.
In the next sections we only use $\mathrm{FPWP}$, given that it is both accurate and much faster than the brute-force method and $\mathrm{PWP}$.
Comparison with alternative methods
In this section we show that both weighting (that is, treating the probabilities as weights) and thresholding (that is, considering the network as a deterministic one after keeping only high-probability edges) lead to wrong estimations of expected modularity. On the contrary, sampling can be used to get more and more accurate results by increasing execution time, so we also study how the balance between time and accuracy when using a sampling approach compares to our algorithm.
Weighting
In this experiment, we show that directly regarding probabilistic networks as weighted deterministic networks leads to a wrong expected modularity calculation. In particular, we first generate a base network using stochastic block modeling, where the parameters are $k=3$, $\mathrm{nc}=9$, $p_{in}=0.72$ and $p_{out}=0.12$. From this base network, we generate multiple probabilistic networks by assigning probabilities to its edges. In each probabilistic network we assign the same probability to all the edges: for example, in Figure 6, the network corresponding to the value 0.20 on the $x$ axis has all its edges having probability 0.20.
@@
Figure 6: Weighted and expected modularity calculated on the same networks.
Thresholding
@@
In this experiment, we compare expected modularity (as computed by $\mathrm{FPWP}$) with the modularity obtained after thresholding, showing that the choice of threshold largely influences the final results, which are also not accurate in general.
@@
Figure 7: Mean and expected modularity calculated by $\mathrm{FPWP}$ and using the threshold method.
(a) CCS network
(b) CCS network
(c) CCS network
(d) LCCS network
(e) LCCS network
(f) LCCS network
@@
Figure 8: Running time and expected modularity value calculated using sampling and $\mathrm{FPWP}$ for networks with a more clear (first row) and less clear (second row) modular structure, with different entropies.
@@
Here we generate planted community structure networks with 110 edges and 3 communities. All networks have the same topology, generated by a stochastic block model with $p_{in}=0.72$ and $p_{out}=0.12$. We then generate different probabilistic networks by assigning random probabilities to the edges so that, for different networks, entropy ratio ranges from 0 to 1, and randomly assign those probabilities to the edges. For each network, we compute the modularities obtained using different thresholds (from 0.1 to 1). In Figure 7 we plot their mean (blue circles) and standard deviation (thick vertical lines). We can see that except for networks with low entropy ratios (that is, networks very close to deterministic networks), the modularities obtained through thresholding are far from the correct values (red line, calculated using $\mathrm{FPWP}$).
Sampling
@@
Thanks to the fact that using $\mathrm{FPWP}$ we can compute expected modularity in polynomial time for a given input network, we can now test the ability of the sampling method to accurately estimate expected modularity, and study the factors influencing the balance between accuracy and execution time when sampling is used.
@@
We use two kinds of random networks. One is characterized by a pronounced community structure, featuring dense intra-community edges and sparse inter-community edges (CCS network). The other exhibits a less distinct community structure, with less intra-community edges and more inter-community edges (LCCS network). Both networks contain 27 nodes and 3 communities. The CCS network and LCCS network are constructed with stochastic block models where the parameters are, respectively, $k=3$, $\mathrm{nc}=9$, $p_{in}=0.99$, and $p_{out}=0.01$, and $k=3$, $\mathrm{nc}=9$, $p_{in}=0.72$ and $p_{out}=0.12$. Notice that the CSS network closely approximates three cliques. When the entropy ratio equals 1.00, all edge probabilities are set to 0.50; when the entropy ratio equals 0.47, all edge probabilities are set to 0.90; when the entropy ratio is 0.00, all edges are deterministic.
@@
Figure 8 shows the execution time (vertical blue line) and the computed expected modularity (blue horizontal line) for our algorithm. Each red circle is the result of a different execution of the sampling method. Specifically, the red nodes arranged in bands’ are the result of different executions of the sampling method, all employing the same input $\theta$ as defined in Algorithm 1. For different executions, we have let the sampling method run for specific amounts of time (starting from the time needed by our algorithm), so that we could inspect the balance between execution time and likelihood to return an accurate result for the sampling method. For example, in Figure 8(f) we can see eight executions of the sampling method close to the blue vertical line: these were stopped after the same time used by our algorithm to return, and their $y$ value is the expected modularity they returned. From the figure we can see that, for different edge probability entropies, sampling always converges to the result produced by $\mathrm{FPWP}$. Except for the case with no uncertainty (entropy ratio=0), sampling takes longer than $\mathrm{FPWP}$ to converge toward the right results. By looking at Figures 8(a)8(b)8(c) or Figures 8(d)8(e)8(f), we find that when entropy ratio increases (larger uncertainty), sampling converges more slowly. If we compare two figures on the same column, e.g. 8(b) and 8(e), we find that when the community structure is less clear, the convergence time of sampling is longer.
Factors influencing execution time of $\mathrm{FPWP}$